# AuscultaBase: A Foundational Step Towards AI-Powered Body Sound Diagnostics

### v1

@@ 6.4 Metric
Macro-F1 and Micro-F1 are reported respectively for each classification task to evaluate the performance across classes and samples. We report both of them because the Macro-F1 scores reflect the model’s ability to handle classes equally well, irrespective of class size, while the Micro-F1 scores emphasize overall classification accuracy across all classes. For the regression task, the Accuracy is computed. Additionally, AUROC (Area Under the Receiver Operating Characteristic) is calculated for binary classification tasks and Class-wise F1 scores for each class are reported for multi-class classification tasks. For a comprehensive overall evaluation, we also report the MRR (Mean Reciprocal Rank) scores and BSC (Borda Count Score) of performance categorized by different perspectives.
7 Results
7.1 Comprehensive Results Overview
The Macro-F1 and Micro-F1 scores for abnormality detection (T1-T8) and disease diagnosis (T9-T15) tasks are exhibited in Table 4 and 5, while the accuracy for the activity detection (T16) task is shown in Figure 2.
Consistent Superiority.
The results in Table 4, 5, and Figure 2 demonstrate that our proposed AuscultaBase-Model outperforms baseline models across various tasks. Specifically, our model attains the top performance on 10 tasks for Macro-F1 and 11 tasks for Micro-F1 out of 15 classification tasks. In contrast, other baselines only achieve superior results on individual tasks. For instance, PANN delivers the highest score only once for the Micro-F1 score (T6) but never achieves the best for the Macro-F1 score. Notably, our model consistently achieves higher performance across various body sounds like lung, heart, and bowel sounds, indicating its superior capability in capturing distinctive and generalizable acoustic features for body sounds.
Balance in Performance.
AuscultaBase-Model achieves a more balanced performance in both Macro-F1 and Micro-F1 scores, which is crucial for classification tasks, especially those involving imbalanced datasets. To evaluate the model’s balanced performance across both metrics, we focus on tasks where the best results were achieved simultaneously for both Macro-F1 and Micro-F1. We find that only 11 tasks (T1, T3-T5, T7-T8, T11-T15) have a model that reaches optimal performance in both metrics. Among these, our model achieves the best results in 9 tasks, while OPERA-CT and CLAP each lead in one task. This balanced optimization is crucial in practical applications, particularly in domains involving imbalanced datasets, where strong performance across all metrics is essential to ensure both minority and majority classes are well-represented.
Metric Function # OPERA-CT AudioMAE CLAP PANN AuscultaBase-Model
Macro-F1 Abnormality Detection 8 0.4688 0.2854 0.6042 0.2375 0.6875
Disease Diagnosis 7 0.3571 0.2571 0.4381 0.3024 0.9286
Overall 15 0.4167 0.2722 0.5267 0.2678 0.8000
Micro-F1 Abnormality Detection 8 0.4583 0.2854 0.4292 0.3396 0.7708
Disease Diagnosis 7 0.3426 0.4500 0.3262 0.3071 0.8571
Overall 15 0.4044 0.3622 0.3811 0.3244 0.8111
Accuracy Activity Detection 1 0.3333 0.2000 0.2500 0.5000 1.0000
Table 6: Mean reciprocal ranks (MRR) grouped by functions and evaluation metrics. The best model within each group is in bold. # means the number of tasks of each group.
7.2 Catergorized Comparison
To provide a broader understanding of the models’ overall effectiveness across various tasks and domains, we categorize these tasks from different perspectives: function, sound type, and task type. Then we adopt MRR (Mean Reciprocal Ranks) and BCS (Borda count scores) metrics of the reported Macro-F1 scores to conduct a comprehensive analysis for them.
Function.
In the function perspective analysis, we compute the MRR across tasks. As shown in Table 6, AuscultaBase-Model consistently achieves the highest MRR in both abnormality detection and disease diagnosis tasks, significantly outperforming other baseline models. Specifically, in disease diagnosis, our model achieved an MRR of 0.9286 for Macro-F1, whereas other models lagged behind, highlighting our model’s capability to not only detect general abnormalities but also make precise and nuanced clinical assessments. This distinction is critical in practical medical applications where the identification of specific diseases can directly impact patient care and treatment decisions. The superior performance of our model across these functions signifies its robustness and precision in both the detection of abnormalities and accurate disease classification, making it particularly valuable for healthcare diagnostics.
Sound Type.
In the sound type analysis, we evaluate the performance of each model using the BCS of their Macro-F1 scores for lung sound, heart sound, and bowel sound tasks. As demonstrated in Figure 3, our AuscultaBase-Model ranks the highest overall, especially excelling in bowel sound tasks. As shown in the left panel of the figure, our model consistently achieves the best performance across different sound types, demonstrating its ability to handle diverse and complex acoustic environments. This strong adaptability makes our model particularly effective for a wide range of body sound classification scenarios.
Task Type.
For the analysis of task type, which includes binary classification, multi-class classification, multi-label classification, and regression, our model also performs exceptionally well. The right panel of Figure 3 shows that our model consistently achieves top performance across all task types especially for binary classification tasks, suggesting a high degree of flexibility and robustness. This is especially important for ensuring that the model can effectively generalize to different problem formulations, ranging from straightforward binary classification to complex multi-label scenarios and regression.
Figure 3: Borda count scores (BCS) scores categorized by the sound type (left) and task type (right).
7.3 AUROC for Binary Classification
To further assess the discriminative power of each model in binary classification tasks, we calculate the AUROC for the binary classification tasks (T3, T13, T14, and T15). Unlike metrics like the F1 score, AUROC provides a more comprehensive evaluation by measuring the model’s ability to distinguish between positive and negative classes at different thresholds.
As shown in Table 7, our model consistently outperforms the other baselines across all tasks, demonstrating its superior capability to distinguish between healthy and abnormal states effectively. For instance, in T13 (heart sound classification), our model achieves an AUROC of 0.917, which is notably higher compared to the best baseline OPERA-CT (0.875). These results underscore the robustness of our model when dealing with challenging binary classification problems, especially those involving subtle variations in body sounds. The results show that our model can effectively capture the critical features necessary for distinguishment, providing a reliable tool for medical diagnostic applications. This consistent superiority across multiple binary classification tasks highlights the overall generalizability and effectiveness of our model in real-world healthcare scenarios.
ID Sound Type OPERA-CT AudioMAE CLAP PANN AuscultaBase-Model
T3 L 0.643$\pm$0.004 0.573$\pm$0.002 0.643$\pm$0.001 0.507$\pm$0.000 0.645$\pm$0.005
T13 H 0.875$\pm$0.001 0.872$\pm$0.004 0.859$\pm$0.001 0.831$\pm$0.002 0.917$\pm$0.003
T14 H 0.606$\pm$0.018 0.635$\pm$0.003 0.628$\pm$0.007 0.540$\pm$0.002 0.659$\pm$0.020
T15 H 0.790$\pm$0.007 0.694$\pm$0.022 0.776$\pm$0.336 0.576$\pm$0.004 0.802$\pm$0.006
Table 7: The AUROC of binary classification tasks (T3, T13-T15). The best results for each task across different auscultation models are in bold.
7.4 Class-wise Performance for Multi-class Tasks
Figure 4: Class-wise F1 scores for multi-class classification tasks (T7, T11, T12). Each label at the corner of the radar chart represents a category, with the number in parentheses indicating the count of samples of that category within the test set. We exclude classes “Bronchi” and “Wheeze&Crackle” for T7 because they are not represented in the test set.
ID # Method OPERA-CT AudioMAE CLAP PANN AuscultaBase-Model
T13 2918 Linear 0.875$\pm$0.001 0.872$\pm$0.004 0.859$\pm$0.001 0.831$\pm$0.002 0.917$\pm$0.003
T15 3294 Linear 0.790$\pm$0.007 0.694$\pm$0.022 0.776$\pm$0.336 0.576$\pm$0.004 0.802$\pm$0.006
Table 8: AUROC for linear probing and full fine-tuning on T13 and T15. # represents the number of training samples for the task.
ID # Metric Method OPERA-CT AudioMAE CLAP PANN AuscultaBase-Model
T11 454 Macro-F1 Linear 31.46$\pm$5.22 11.83$\pm$0.00 31.69$\pm$4.05 31.45$\pm$6.85 41.54 $\pm$3.44
Micro-F1 Linear 38.00$\pm$2.00 42.00$\pm$2.00 45.00$\pm$1.00 43.00$\pm$3.00 50.00$\pm$2.00
Table 9: Macro-F1 and Micro-F1 scores for linear probing and full finetuning on T11.
Building upon the previous evaluations which present a comprehensive analysis, this section aims to delve deeper into how each model handles individual classes, especially under imbalanced conditions. To this end, we plot the F1 scores of different models on multi-class classification tasks (T7, T11, T12) using radar charts. To make a clearer comparison, the results for each class are applied with Min-Max normalization across different models. As shown in Figure 4, our model (depicted in red) consistently shows superior or competitive F1 scores across both majority and minority classes.
For example, in T7 (which is highly imbalanced as shown in Table 2), our model achieves higher scores for the minority categories such as “Crepitation” and “Wheeze”, indicating its capability to effectively learn meaningful features even from underrepresented classes. This balanced performance across categories, irrespective of their distribution, highlights our model’s robustness and effectiveness in dealing with skewed datasets.
For T11, which involves classifying the severity of COPD, the class distribution is also significantly imbalanced, with “COPD4” having a much higher number of samples compared to “COPD0” and “COPD1”. Despite this challenge, Figure 4 indicates that our model still achieves high F1 scores not only in the majority class (“COPD4”) but also in minority classes like “COPD0”, suggesting that our model can handle class imbalance effectively without overly favoring majority classes. This indicates a good generalization capability and adaptability in identifying underrepresented classes.
Similarly, for T12, although the overall and test set distributions are more balanced across classes (e.g., each class has around 200 samples and shares a balanced number of test samples), our model still achieves superior F1 scores in multiple categories. This suggests that our model not only excels in imbalanced scenarios but also maintains superiority when the class distribution is relatively uniform.
7.5 Full Fine-tuning Performance
In this section, we highlight the performance of different models under full finetuning, focusing on both binary and multi-class classification tasks. For binary tasks such as T13 and T15, Table 8 depicts that our model consistently achieves the highest AUROC scores after full finetuning. Specifically, on T13, our model reached an AUROC of 0.978, which makes it highly effective for clinical applications that require accurate identification of specific conditions.
For the multi-class classification task T11 as shown in Table 9, which features a relatively small number of training samples (454), our model demonstrates notable improvements after full finetuning, achieving a Macro-F1 of 55.14 and a Micro-F1 of 57.00. The performance gain over linear finetuning suggests that our model is highly adaptable, effectively capturing complex class relationships even when training data is limited. This adaptability is crucial for real-world medical diagnostics, where imbalanced datasets and limited data availability are common challenges. Our model’s ability to excel under such conditions further underscores its robustness and utility in various healthcare scenarios.
8 Conclusion
This work presents a large-scale body sound corpus (A4Body-Corpus), an auscultation acoustic foundation model (A4Body-Model), and a comprehensive benchmark for evaluation (A4Body-Bench). Our work addresses existing limitations in traditional auscultation practices through the integration of AI. By training on a diverse collection of body sounds using contrastive learning, our model shows a substantial improvement over existing pre-trained acoustic models, especially in imbalanced and limited-data settings. The full fine-tuning results reveal the robustness of our model in both binary and multi-class classification tasks, underscoring its potential for real-world clinical applications. This work emphasizes the value of combining traditional auscultation with AI-driven methods to deliver more reliable and accessible healthcare solutions, paving the way for the broader adoption of AI in medical diagnostics.
Appendix A Datasets
A.1 Datasets Availability
@@
We utilized 11 datasets in our benchmark, with access methods and licensing details provided in Table 10. A more detailed description of each dataset is provided below. Notably, all datasets include both audio recordings and accompanying metadata. The audio data has been anonymized, and the metadata contains no personally identifiable information or offensive content.
@@ A.2 Datasets Description
This dataset consists of 8, 377 auscultation audios recorded by 8, 377 children participants without diagnostic labels. The auscultation devices, positions, and procedures are followed with those of HSDReport. This research is carried out with a strong focus on ethics, protecting participant privacy, and following all rules. Before taking part, all individuals or their legal guardians were given full details about the study’s goals, what data would be collected (such as heart sounds and ultrasound results), and how this data would be used for research. Consent was obtained only after making sure everyone understood these details. This dataset is collected by Xinhua Hospital Affiliated To Shanghai Jiao Tong University School of Medicine, and due to the copyright constraint, it has not been open source yet.
@@
Appendix B Demographic and Clinical Information
B.1 Age and Gender
@@ B.3 Sound Spectrogram
Appendix C Downstream Task Description
@@
  • Binary Classification (Tasks 3, 13, 14, 15): Tasks that require predicting a binary outcome (normal/abnormal) from an audio recording.
  • Multi-Class Classification (Tasks 1-2, 4-9, 11-12): The tasks involve classifying an audio recording into one or multiple classes of several predefined categories.
  • Multi-Label Classification (Task 10): The task involves classifying an audio recording into one of several predefined categories.
  • Regression (Task 16): The task of predicting continuous values from audio data, such as the count of bowel sounds.
@@
Task 6: LungSound_MC_LungSound.
@@
Task 7: ICBHI2017_MC_LungSound.
Task 8: Circor2022_MC_Murmur.
@@
Task 9: ICBHI2017_MC_Disease.
@@
Task 11: RD@TR_MC_Disease.
Task 11 is a five-class classification task that categorizes the data into COPD severity levels ranging from 0 to 4 based on audio. The goal of this task is to predict the severity of COPD using lung sounds.
Task 12: Korean_MC_Disease.
@@
Task 16: BowelSound_R_Count.
Task 16 is a regression task aimed at assessing the frequency of specific intestinal activities or patterns by counting the occurrence of bowel sounds in the recordings.
Appendix D Implementation Details
D.1 Pretraining
Data Processing
We pre-train our model using a combination of 11 datasets as described in Table 1. We apply validation with the validation sets of the datasets (e.g. SPRSound, HF Lung, ICBHI 2017, and Bowel Sound). For those datasets that do not have an existing validation set, we randomly split 10% of the data to serve for validation. To adapt to the varied lengths of recordings, we adopt random cropping for each audio and feed them to the pertaining. according to the minimum duration of the recordings for each dataset, different crop lengths are set. Specifically, the cropped length for Korean and Bowel Sound is set as 320 ms, while it is 640 ms for others. Additionally, as a usual augmentation technique, we adjust the loudness of the recording (Schlüter & Grill, 2015) with a random scale factor ranging from 0.9 to 1.1. Moreover, SpecAugment (Park et al., 2019) is also adopted.
Constrastive Learning-based Method
The target of our method is to guide the model to distinguish the source of audio segments, where the two segments from the same spectrogram are regarded as a positive pair, otherwise negative pairs. Given two segments $x$ and $x^{\prime}$, the features will first be extracted with an encoder $f(\cdot)$ from them, and then a projector $g(\cdot)$ will map the features into a lower dimensional space to obtain the final representation. To improve the quality and discriminative power of the learned representations, an extra linear layer $W$ is added on top of the whole model for the positive/negative anchor $x^{\prime}$. Then the bilinear similarity $s(\cdot)$ is calculated as
$s(x,x^{\prime})=g(f(x))^{T}Wg(f(x^{\prime})).$ (1)
D.2 Benchmarking
Data Processing
In the evaluation stage, we adopt the validation set of the pretraining stage as the test set, which is always held out for pretraining and downstream fine-tuning. Considering the varied lengths of different recordings even within the same dataset, each audio in the training set is chunked and padded into the same length. Specifically, each recording of all the datasets is chunked into 8s except for T14 (Korean dataset) and T18 (Bowel Sound dataset). We choose 8s to satisfy the input requirements of different models as shown in xx. As for the Korean and Bowel Sound datasets, the maximum durations are set as 4s and 2s respectively because their recordings are rather short as demonstrated in Table 1.
Baselines
@@ Appendix F Additional Evaluation Results
F.1 Performance Radar Plot
Figure 12: Radar plot of reciprocal ranks of Macro-F1 and Micro-F1 scores for classification tasks respectively.
Apart from the mean reciprocal ranks tasks depicted in Table 6, we demonstrate the reciprocal ranks of all the classification tasks (T3-T17) which is detailed in Figure 12.

### v2

@@ 4.3.5 Evaluation Metric
For performance evaluation, Macro-F1 and Micro-F1 are reported respectively for each classification task to evaluate the performance across classes and samples. We report both of them because the Macro-F1 scores reflect the model’s ability to handle classes equally well, irrespective of class size, while the Micro-F1 scores emphasize overall classification accuracy across all classes. Additionally, Precision and Recall scores are also reported for each task. The ROC (Receiver Operating Characteristic) curve and AUC (Area Under the ROC curve) are also plotted and calculated for binary classification tasks. For a comprehensive overall evaluation, we also report the BCS (Borda Count Score) of performance categorized by different perspectives.
For comparative analysis between our model and human experts, we compared the diagnostic outcomes using standard metrics: confusion matrices, sensitivity, specificity, and overall accuracy. The confusion matrix provides a detailed breakdown of true positives (TP), false positives (FP), false negatives (FN), and true negatives (TN). Sensitivity reflects the ability to correctly identify diseased cases, while specificity indicates the correct identification of healthy cases. Accuracy, the overall proportion of correct predictions, provides a holistic measure of diagnostic performance. These metrics are formulated as:
$Sensitivity=\frac{TP}{TP+FN},Specificity=\frac{TN}{FP+TN},Accuracy=\frac{TP+TN}{TP+FP+FN+TN}.$ (3)
Ethical Considerations
This research was conducted in accordance with established ethical principles for medical studies, ensuring the rights, privacy, and well-being of all participants. Below is an overview of the primary ethical considerations addressed.
Informed Consent
Prior to any data collection, all participants or their legal representatives received detailed information about the objectives, methods, and potential applications of this study. Written informed consent was obtained from every participant, aligning with international standards such as the Declaration of Helsinki. This process guaranteed that individuals were fully aware of both the scope and implications of their involvement in the research.
Data Confidentiality and Security
To protect participant privacy, strict measures were implemented to remove any personal identifiers from collected data, ensuring anonymity. Additionally, encrypted storage systems with controlled access were employed to maintain data security and prevent unauthorized viewing or manipulation. These steps were critical in upholding the confidentiality of the information throughout the study.
Ethical Review and Oversight
This study protocol underwent a comprehensive evaluation by an independent ethics committee, receiving approval under the reference number XHEC-C-2020-009-1. The committee provided ongoing oversight to confirm adherence to ethical standards and safeguard participant welfare at every stage of the project.
Data Availability
@@
We utilized 11 datasets in our benchmark, with access methods and licensing details provided in Extended Table 5. Notably, most of datasets include both audio recordings and accompanying metadata. The audio data has been anonymized, and the metadata contains no personally identifiable information or offensive content.
Code Availability
The code and model checkpoint has been released in https://github.com/applewpj/AuscultaBase.
Appendix A Related Works
The auscultatory body sounds encompass heart, respiratory, and bowel sounds, with most existing models focusing on a single category. For heart sounds, the primary tasks include abnormality detection and disease diagnosis. Traditional methods often rely on acoustic features such as STFT for feature processing. For instance, Chen et al. (2023) applied noise reduction followed by STFT to obtain spectrograms, using a CNN with an attention module for classification. Similarly, Guo et al. (2023) enhanced feature extraction by combining high-order spectral estimation with STFT, employing a dual-stream CNN as the primary architecture. General audio pre-training models have also been explored. For example, (Koike et al., 2020) compared PANN-extracted features (Kong et al., 2020) with STFT-derived features used in non-pre-trained architectures like VGG and ResNet, demonstrating the transferable capabilities of pre-trained models in heart sound diagnosis.
Respiratory sound analysis similarly focuses on abnormality detection and disease diagnosis. Early approaches utilized traditional feature extraction techniques such as STFT and MFCC (Oletic & Bilas, 2017; Jung et al., 2021). Recently, the COVID-19 pandemic spurred the development of pre-training models for respiratory sounds. Zhang et al. (2024) constructed a dataset combining cough and stethoscope-collected respiratory sounds, comparing contrastive and generative pre-training methods across 19 tasks. Additionally, Baur et al. (2024) developed an acoustic event detector to identify respiratory-related events from online audio sources, leveraging a generative pre-training approach based on MAE (He et al., 2022).
Bowel sounds remain underexplored due to a lack of datasets and established methods. Current work focuses primarily on bowel sound recognition. For instance, Sitaula et al. (2022) collected data from 49 newborns and developed a hybrid CNN and hidden semi-Markov model using MFCC features. Similarly, Baronetto et al. (2023) investigated the utility of AudioSet (Gemmeke et al., 2017) for pre-training and evaluated CNN and attention-based architectures on data from 18 participants.
While these efforts represent progress, they fail to address a critical gap: the intrinsic inter-relatedness of body sounds in auscultation. Existing methods operate in silos, treating each body sound category independently and overlooking shared patterns and synergies. Furthermore, no existing approach attempts to create a unified framework capable of leveraging the collective knowledge of all auscultation data.
To overcome these limitations, we introduce AuscultaBase, the first-ever audio foundation model designed specifically for body sound auscultation. Inspired by the transformative success of foundation models in vision and language domains, AuscultaBase introduces a multi-layered knowledge integration paradigm that unifies data across heart, respiratory, and bowel sound categories. It achieves this through an adaptive acoustic fusion mechanism, enabling dynamic representation learning from diverse audio sources.
Appendix B Datasets
B.1 Datasets Availability
@@ B.2 Datasets Description
This dataset consists of 8, 377 auscultation audios recorded by 8, 377 children participants without diagnostic labels. The auscultation devices, positions, and procedures are followed with those of HSDReport. This research is carried out with a strong focus on ethics, protecting participant privacy, and following all rules. Before taking part, all individuals or their legal guardians were given full details about the study’s goals, what data would be collected (such as heart sounds and ultrasound results), and how this data would be used for research. Consent was obtained only after making sure everyone understood these details. This dataset is collected by Anonymous Hospital, and due to the copyright constraint, it has not been open source yet.
@@
Appendix C Demographic and Clinical Information
C.1 Age and Gender
@@ C.3 Sound Spectrogram
Appendix D Downstream Task Description
@@
  • Binary Classification (Tasks 3, 9, 14, 15, 16): Tasks that require predicting a binary outcome (normal/abnormal) from an audio recording.
  • Multi-Class Classification (Tasks 1-2, 4-8, 10, 12-13): The tasks involve classifying an audio recording into one or multiple classes of several predefined categories.
  • Multi-Label Classification (Task 11): The task involves classifying an audio recording into one of several predefined categories.
@@
Task 6: ICBHI2017_MC.
Task 7: LungSound_MC.
@@
Task 8: Circor2022_MC.
@@
Task 9: BowelSound_BC.
Task 9 is a binary classification task aimed at assessing the occurrence of bowel sounds in the recordings.
Task 10: ICBHI2017_MC_Disease.
@@
Task 12: RD@TR_MC.
Task 11 is a five-class classification task that categorizes the data into COPD severity levels ranging from 0 (under risk) to 4 (very severe level) based on audio. For a more detailed description of the grading criteria, please refer to Altan et al. (2017) and Roisin (2016). The goal of this task is to predict the severity of COPD using lung sounds.
Task 13: Korean_MC.
@@ Appendix F Additional Evaluation Results
F.1 Detailed Class-wise Evaluation in Multi-class Scenarios
Figure 13: Class-wise F1 scores for multi-class classification tasks (T7, T12, T13). Each label at the corner of the radar chart represents a category, with the number in parentheses indicating the count of samples of that category within the test set. We exclude classes “Bronchi” and “Wheeze & Crackle” for T7 because they are not represented in the test set.
Figure 13 presents radar charts of class-wise F1 scores for multi-class classification tasks (T7, T12, and T13). Our model (highlighted in red) consistently demonstrates competitive or superior performance across both majority and minority classes. For instance, in T7—a task marked by a significant class imbalance—AuscultaBase outperforms competitors in minority categories such as “Crepitation” and “Wheeze”. Similarly, in T12, which involves grading the severity of COPD (ranging from COPD0 to COPD4), our model maintains high F1 scores even for the underrepresented classes. This balanced performance across classes underlines the model’s robustness in handling real-world datasets, where class imbalances are common and can lead to biased diagnostic outcomes.
F.2 Enhanced Performance Under Full Fine-tuning
We further investigate model performance under full fine-tuning conditions for both binary and multi-class tasks. For binary tasks (T14 and T16), Table 7 shows that AuscultaBase achieves significantly improved AUC scores after full fine-tuning, with task T14 reaching an AUC of 0.978. Such improvements are particularly valuable in clinical scenarios that demand high sensitivity and specificity. For the multi-class task T12—despite its relatively small training sample size (454 samples)—AuscultaBase achieved a Macro-F1 of 41.78 (linear) and 55.14 (full fine-tuning), with corresponding Micro-F1 scores of 49.00 and 57.00, respectively (Table 8). These gains underscore the model’s adaptability and its capacity to capture complex inter-class relationships, even when data are limited—a common challenge in medical diagnostics.
ID # Method OPERA-CT AudioMAE CLAP PANN AuscultaBase
T14 2918 Linear 0.850$\pm$0.003 0.872$\pm$0.003 0.835$\pm$0.023 0.824$\pm$0.001 0.920$\pm$0.008
T16 3294 Linear 0.795$\pm$0.005 0.691$\pm$0.022 0.766$\pm$0.010 0.484$\pm$0.020 0.802$\pm$0.008
Table 7: AUC for linear probing and full fine-tuning on T14 and T16. # represents the number of training samples for the task. AUC ranges from 0 to 1, and higher is better.
ID # Metric Method OPERA-CT AudioMAE CLAP PANN AuscultaBase
T12 454 Macro-F1 Linear 30.74$\pm$0.77 17.00$\pm$5.17 27.68$\pm$2.07 30.15$\pm$5.57 41.78$\pm$1.83
Micro-F1 Linear 36.00$\pm$2.00 43.00$\pm$1.00 41.00$\pm$3.00 40.00$\pm$4.00 49.00$\pm$1.00
Table 8: Macro-F1 and Micro-F1 scores for linear probing and full finetuning on T12. Macro-F1 and Micro-F1 range from 0 to 100, and higher is better.